OpenStudy (anonymous):

\[\LARGE \int \frac{9-\sqrt x}{3-\sqrt x}dx\] Can't even start... need some help.

OpenStudy (anonymous):

I think maybe writing u = sqrt(x) and du = (1/2)1/sqrt(x) might simplify a bit, but if you follow this way you will have to do long division again... Wonder if there is a smart trick to simplify.

OpenStudy (experimentx):

\[ \frac{9 - \sqrt x}{3 - \sqrt x} = 1 + \frac{6}{3 -\sqrt x}\] I think ...

OpenStudy (anonymous):

^ That looks easier compared to what I got.

OpenStudy (anonymous):

well @bmp I guess you have to answer this first lol http://openstudy.com/study#/updates/4fa03ee2e4b029e9dc314772 @experimentX ...how did that come, I want to know how to do it, I can't keep asking for every integral I have :(

OpenStudy (anonymous):

It's just a way of polynomial division: http://en.wikipedia.org/wiki/Polynomial_long_division

OpenStudy (anonymous):

so, without knowing how do to "long division" I'm not actually capable to solve these integrals , am I ? :(

OpenStudy (anonymous):

Well, if you learned some other way of doing polynomial division, it should work fine. But if don't know that, I guess, yeah, you should look into learning it :-)

OpenStudy (anonymous):

and if try this: \[(6+3-\sqrt{x})/3-\sqrt{x}=6/(3-\sqrt{x}) +1\]

OpenStudy (anonymous):

KhanAcademy has videos to cover that, tho. So, it shouldn't be too hard to pick up.

OpenStudy (anonymous):

later make variable change x=t^2 and should be easy

OpenStudy (anonymous):

ahh that's cool then, I'll just check khan academy. Thanks everyone for help\suggestions and everything... I'm wondering, to who should I give a medal now? :S ... (You give medals to each other, and I'll give one to the last) .. ;)

OpenStudy (anonymous):

No problem. And polynomial division seems daunting at first because it's messy, but it's not that hard :-)

OpenStudy (anonymous):

;) ... I'll see what I can do. ♥